Today volcanic activity was back to a similar level to that before yesterdays pyroclastic flows. Seismicity was dominated by rockfall signals with a few small hybrid earthquakes. Visibility was poor and no clear observations were made of the dome.
Between 16:00 yesterday and 16:00 today there were 21 long-period earthquakes, 23 hybrid earthquakes, and 121 rockfall signals. One volcano-tectonic earthquake was recorded.
A GPS survey was carried out today of the pyroclastic flow deposits in Tuitt's, Mosquito and White's Ghauts. The results from this survey will be reported at a later date.
The pyroclastic flows yesterday show that the dome is still growing. Further flows could occur down any of the ghauts around the volcano with little warning. The level of ash in the atmosphere in Woodlands, St Peter's and in St John's, as averaged over the last 24 hours, has remained low.